OpenStudy (anonymous):

Which of the following is a possible rational zero for the polynomial 3x3 + 5x2 - 4x + 4? A. -3 B. 0 C. 23 D. 34

OpenStudy (anonymous):

i say that its b

OpenStudy (anonymous):

well did you try simplifying the polynomial??

OpenStudy (anonymous):

Why? Can you explain what you did?

OpenStudy (anonymous):

well you have this equation 3x^3 + 5x^2 - 4x + 4? because it is a polynomial you must separate the two. (3x^3 + 5x^2) + (-4x - 4) and then you simplify x^2(3x + 5) -4(x - 1) well then the zeros will be 0, -5/3, 1

OpenStudy (john_es):

None of the answers can be a zero of the polynomial.

OpenStudy (anonymous):

So, @FriedRice how do I determine the zeroes after simplifying? What did you do to the simplified polynomial?

OpenStudy (anonymous):

Why do you say that @John_ES ?

OpenStudy (john_es):

I say that the zeroes of the function, \[3 x^3 + 5 x^2 - 4 x + 4\] is none of the given answers.

OpenStudy (anonymous):

first i separated the polynomial into 2 parts as i have shown above. then i simplified it

OpenStudy (anonymous):

then you have to make one of each x = 0

OpenStudy (anonymous):

Yes I know that, you showed me that you got \[x^2(3x+5)-4(x-1)\]But how did you get the zeroes from that? @FriedRice

OpenStudy (anonymous):

here an example from your question x^2 = 0 so x = 0 (3x + 5) = 0 so x = 5/3 but because you have only -4 you dont have to do anything then (x - 1) = 0 so x = 1

OpenStudy (john_es):

But neither of these numbers are zeroes of the function. You can plug in them, and you won't obtain 0.

OpenStudy (anonymous):

They aren't real zeroes

OpenStudy (anonymous):

They're rational zeroes

OpenStudy (anonymous):

If you graphed this polynomial, you wouldn't see the graph intersect in any of these locations. However, they are still considered zeroes.

OpenStudy (anonymous):

So @FriedRice I need to set each binomial to 0? (3x+5) (x-1) and x^2-4)?

OpenStudy (anonymous):

hmmm then i have a another explaination.well because this is the rational roots theorem; essentially they must have a numerator that divides our constant (here 4) and a denominator which divides our leading coefficient (here 3)... note that 3 has divisors ±1,±3 and 4 has divisors ±1,±2,±4.

OpenStudy (anonymous):

So would 3/4 work?

OpenStudy (anonymous):

so the answer to what i suspect is option c

OpenStudy (john_es):

Well, this is the first time I know about such zeroes. I only know these, http://www.sosmath.com/algebra/factor/fac10/fac10.html

OpenStudy (anonymous):

Oh okay, I mixed up numerator and denominator there, sorry

OpenStudy (anonymous):

Yes @John_ES I've been to that site too, but I don't understand it. I really just need someone to walk me through on of these

OpenStudy (anonymous):

so do you understand now?

OpenStudy (anonymous):

Somewhat. I just need to list the factors of the constant and leading coefficient and then make a fraction using the factors? @FriedRice

OpenStudy (john_es):

As I said before, rational zeros or zeros of the polinomy must be roots of the polinomy. Only to clarify, rational zeroes are roots of the function.

OpenStudy (anonymous):

just listen to what john_es has to say but yes @graham69

OpenStudy (anonymous):

Okay, thank you @FriedRice

OpenStudy (anonymous):

no prob

OpenStudy (john_es):

And of course, the last answer of @FriedRice is the correct one. 2/3

OpenStudy (john_es):

This is a possible zero, but not the true zero of the function.

OpenStudy (anonymous):

Alright, so if I had the polynomial \[4x^3+7x^2-5x+3\]I would take the divisors of three as ±1 and ±3, and the divisors of 4 as ±1 ±2 ±4. Then I just need to make a fraction out of them? @John_ES

OpenStudy (john_es):

In general, if you have a polinomy of the type, \[ax^3+\ldots+c=0\]Then the possible zeros are divisors of c/a (positives and negatives).

OpenStudy (john_es):

Yes, you can do as you said.

OpenStudy (anonymous):

Awesome. Thank you so much @John_ES

OpenStudy (john_es):

You're welcome.